Circular No. 8200

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003hy IN IC 5145 T. Boles, Coddenham, England, reports the discovery of an apparent

supernova (mag 16.6) on unfiltered CCD images taken on Sept. 14.007 and 14.837 UT with a 0.35-m reflector in the course of the U.K. Nova/Supernova Patrol. The new object is located at α = 21^{h}54^{m}22^{s}.72, δ = +15^{o}09′37′′.8, which is ≈ 5′′.5 west and 12′′.5 north of the center of IC 5145. SN 2003hy is not present on Boles’ images from Aug. 16 and Sept. 1 (limiting mag 19.5) or on Palomar Sky Survey red (1991) or blue (1988) plates. M. Armstrong, Rolvenden, England, reports his independent discovery of SN 2003hy at mag 16.3 on unfiltered CCD images taken on Sept. 13.865 and 14.815, also with a 0.35-m reflector in the course of the U.K. Nova/Supernova Patrol; Armstrong provides position end figures 22^{s}.73, 37′′.2. Nothing was visible on Armstrong’s images taken on Aug. 3 (limiting mag 19.5) and Sept. 4 (limiting mag 18.5).

SUPERNOVA 2003hx IN NGC 2076 M. Salvo, J. Norris, and B. Schmidt, Australian National University

(ANU), report that a spectrum (range 380–900 nm) of SN 2003hx (cf. IAUC 8199), taken on Sept. 13.78 UT at the ANU 2.3-m telescope (+ Double Beam Spectrograph), shows it to be a type-Ia supernova around maximum light. The minimum of the Si II 635.5-nm line gives an expansion velocity of ∼ 12000 km/s if the NED recession velocity of 2142 km/s for the host galaxy is adopted.

V475 SCUTI = NOVA SCUTI 2003 N. N. Samus, Institute of Astronomy, Moscow, informs us that the

designation V475 Sct has been given to this nova (cf. IAUC 8190, 8199).

CCD magnitudes: Aug. 30.986 UT, V = 9.15 (D. Rodriguez, Madrid, Spain); 31.017, V = 9.26 (W. Liller, Viña del Mar, Chile; through thin clouds); Sept. 3.078, V = 8.63 (J. D. West, Mulvane, KS); 3.084, R_{c} = 8.07 (West). Visual magnitude estimates: Aug. 31.011 UT, 9.5 (J. G. S. Aguiar, Campinas, Brazil); 31.863, 9.0 (P. Schmeer, Bischmisheim, Germany); Sept. 1.799, 8.7 (V. Nevski, Vitebsk, Belarus); 2.860, 8.6 (A. Baransky, Bucha, Ukraine); 3.867, 8.9 (A. Diepvens, Balen, Belgium); 5.909, 8.9 (K. Hornoch, Lelekovice, Czech Republic); 6.771, 8.8 (S. E. Shurpakov, Baran’, Belarus); 7.847, 9.2 (Baransky); 9.806, 8.6 (Shurpakov); 13.937, 9.1 (E. Broens, Mol, Belgium); 14.835, 9.1 (Broens).

2003 September 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green